Jackson Company is considering a project that has the following cash flows. What are the project's payback, discounted payback, NPV, IRR, and MIRR? The weighted Average Cost of capital of Jackson Company is 10 percent. Explain, in writing, if the project is accepted and why?

Year

Cash Flows ($)

0

-950

1

525

2

485

3

445

4

405
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

A В C Е F G Н 10% Cost of capital 2. 3 PVF(10%) 4 Cash Flow Cumulative CF PV CF Cumulative PV CF Year (950.00) 5 (950.00) (95

Cell reference -

А В C E F Cost of capital 0.1 2 PVF(10% =1/(1+$C$2)^B5C5*E5F5 = 1/(1+$C$2)^B6C6*E6G5+F6 = 1/(1+$C$2)^B7C7*E7G6+F7 = 1/(1+$C$2

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.